OpenStudy (anonymous):

In partial fraction,can we factorize the numerator(2x+8/(x+5)(x+3))?

OpenStudy (lgbasallote):

uhmm you can..but it wouldnt make any sense? lol

OpenStudy (anonymous):

no,we factorize denominator alwayssssssss

OpenStudy (anonymous):

what so you mean? like my teacher said before we begin we need to factorise.but i don't know which one

OpenStudy (anonymous):

*do

OpenStudy (anonymous):

@nitz Always denominator?

OpenStudy (anonymous):

see the resolution of f(x)/g(x) into partial fractions depends mainly upon the nature of the factors of g(x)

OpenStudy (anonymous):

ie the denominator,,,,,

OpenStudy (anonymous):

oh!!! ok..ok I see @nitz can you solve this for me pls.

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

\[A/(X+5) + B/(X+3) = (2X+8)/(X+5)(X+3) \]

OpenStudy (anonymous):

2x+8/(x+5)(x+3)=A/(x+5)+B/(x+3) 2x+8/(x+5)(x+3)=(A(x+3)+B(x+5))/(x+5)(x+3)

OpenStudy (anonymous):

cancel the common factors ie (x+3)(x+5) on both sides assuming them to be non zero

OpenStudy (anonymous):

we are left with 2x+8=A(x+3)+B(x+5)

OpenStudy (anonymous):

put x=-3,we get 2(-3)+8=A(-3+3)+B(-3+5) 2=2B B=1

OpenStudy (anonymous):

SImilarly put x=-5 YOU WILL GET A =1

OpenStudy (anonymous):